# PDRL: Post-hoc Descriptor-based Residual Learning for Uncertainty-Aware Machine Learning Potentials

### v1

@@ Appendix C Training details and hyperparameters
During the PDRL experiments, validation set was used for learning rate scheduling and early stopping. Training began with a learning rate of and patience of 10 epochs, halving the learning rate whenever the validation error failed to improve for 10 consecutive epochs. Training was terminated after a maximum of 1000 epochs or once the learning rate decreased to .
[…]
@@
Appendix D Additional experimental results
D.1 AUC evaluation of in-domain dataset
@@
[…]
| MC-dropout [Gal and Ghahramani, 2016] | -0.48 (0.01) | 0.49 (0.02) | 0.50 (0.01) | 0.49 (0.04) | 0.55 (0.02) | |
| GMM [Zhu et al., 2023] | 0.47 (0.06) | 0.46 (0.07) | 0.55 (0.04) | 0.71 (0.04) | 0.54 (0.01) | |
| kNN [Janet et al., 2019] | 0.48 (0.05) | 0.47 (0.05) | 0.54 (0.03) | 0.70 (0.04) | 0.51 (0.01) | |
| PDRL-norm | 0.51 (0.03) | 0.48 (0.01) | 0.46 (0.02) | 0.72 (0.03) | 0.63 (0.02) | |
| PDRL-diff | 0.56 (0.07) | 0.57 (0.06) | 0.59 (0.06) | 0.90 (0.04) | 0.61 (0.01) | |
| MC-dropout [Gal and Ghahramani, 2016] | 0.62 (0.03) | 0.63 (0.02) | 0.63 (0.04) | 0.95 (0.01) | 0.83 (0.01) | |
| GMM [Zhu et al., 2023] | 0.79 (0.00) | 0.84 (0.02) | 0.82 (0.02) | 0.98 (0.01) | 0.85 (0.03) | |
| kNN [Janet et al., 2019] | 0.77 (0.01) | 0.81 (0.02) | 0.81 (0.02) | 0.98 (0.01) | 0.84 (0.00) | |
| PDRL-norm | 0.85 (0.01) | 0.86 (0.03) | 0.83 (0.01) | 0.98 (0.00) | 0.98 (0.00) | |
| PDRL-diff | 0.76 (0.01) | 0.78 (0.02) | 0.77 (0.02) | 0.96 (0.06) | 0.92 (0.01) |
D.2 OOD detection results with standard deviation
@@
| MC-dropout [Gal and Ghahramani, 2016] | 0.92 (0.01) | 0.54 (0.10) | 0.81 (0.03) | 0.62 (0.06) | 0.72 |
| GMM [Zhu et al., 2023] | 0.98 (0.00) | 0.74 (0.05) | 0.73(0.02) | 0.77 (0.04) | 0.81 |
| kNN [Janet et al., 2019] | 0.98 (0.01) | 0.80 (0.03) | 0.75 (0.03) | 0.76 (0.07) | 0.82 |
| PDRL-norm | 0.99 (0.00) | 0.82 (0.10) | 0.78 (0.26) | 0.70 (0.06) | 0.82 |
| PDRL-diff | 0.97 (0.00) | 0.89 (0.04) | 0.80 (0.08) | 0.81 (0.03) | 0.87 |
@@
| MC-dropout [Gal and Ghahramani, 2016] | 1.00 (0.00) | 0.63 (0.04) | 0.84 (0.04) | 0.82 (0.08) | 0.82 |
| GMM [Zhu et al., 2023] | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | 1.00 |
| kNN [Janet et al., 2019] | 1.00 (0.00) | 0.99 (0.01) | 1.00 (0.00) | 1.00 (0.00) | 1.00 |
| PDRL-norm | 1.00 (0.00) | 0.82 (0.09) | 0.82 (0.25) | 0.99 (0.02) | 0.91 |
| PDRL-diff | 1.00 (0.00) | 0.97 (0.04) | 1.00 (0.00) | 1.00 (0.00) | 0.99 |
Appendix E When does PDRL outperform kNN and GMM?: an analysis based on principle component analysis (PCA)
[…]
Similar trends are observed in Figure 3, which only shows the calcium atoms in HME21. The top-left region of the PCA plot is densely populated, yet exhibits relatively high force-prediction errors. Both of our PDRL methods successfully capture this behavior, as the prediction error was explicitly incorporated into the training process. In contrast, the kNN and GMM approaches rely solely on descriptor information and therefore incorrectly assign low uncertainty to the same tail region in the top left. These results suggest that PDRL methods have strong potential for uncertainty prediction in more diverse and complex datasets, paving the way toward the development of universal potentials uncertainty estimation.
@@ Appendix E When does PDRL outperform kNN and GMM?: an analysis based on principle component analysis (PCA)
Justification: We propose PDRL and evaluate its performance to show the capability of our method.
@@
Justification: We share the training information and training parameters in appendix and disclose training hyperparameters in appendix.
@@
Justification: We do not plan to release the code at this moment.
@@
[…]
@@
Justification: We described this information in Appendix C.
@@
[…]

### v2

@@ Appendix B Training details and hyperparameters
During the P-DRUM experiments, validation set was used for learning rate scheduling and early stopping. Training began with a learning rate of and patience of 10 epochs, halving the learning rate whenever the validation error failed to improve for 10 consecutive epochs. Training was terminated after a maximum of 1000 epochs or once the learning rate decreased to .
Except for and HME21 dataset, in which we used a batch size of for P-DRUM-diff forces prediction, the batch size was set to atoms in all the other P-DRUM forces training. For P-DRUM energy training, the batch size was set to chemical structures. The larger dataset sizes in these cases lead to a greater number of total atoms per batch, and increasing the batch size helped stabilize training. For P-DRUM-norm energy and forces prediction, we employed the ReLU activation function with a single hidden layer, along with a softplus activation right before output. P-DRUM-diff uses similar but softplus-removed MLP architecture, where one hidden layer was used for energy and two hidden layers were used for forces. For MC-dropout, we set the dropout ratio to 10%.
@@
Appendix C Additional experimental results
C.1 AUC evaluation of in-domain dataset
@@
C.2 OOD detection results with standard deviation
@@
| MC-dropout [gal2016dropout] | 0.92 (0.01) | 0.54 (0.10) | 0.81 (0.03) | 0.62 (0.06) | 0.72 |
| GMM [zhu2023fast] | 0.98 (0.00) | 0.74 (0.05) | 0.73(0.02) | 0.77 (0.04) | 0.81 |
| kNN [janet2019quantitative] | 0.98 (0.01) | 0.80 (0.03) | 0.75 (0.03) | 0.76 (0.07) | 0.82 |
| P-DRUM-norm | 0.99 (0.00) | 0.82 (0.10) | 0.78 (0.26) | 0.70 (0.06) | 0.82 |
| P-DRUM-diff | 0.97 (0.00) | 0.89 (0.04) | 0.80 (0.08) | 0.81 (0.03) | 0.87 |
@@
| MC-dropout [gal2016dropout] | 1.00 (0.00) | 0.63 (0.04) | 0.84 (0.04) | 0.82 (0.08) | 0.82 |
| GMM [zhu2023fast] | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | 1.00 |
| kNN [janet2019quantitative] | 1.00 (0.00) | 0.99 (0.01) | 1.00 (0.00) | 1.00 (0.00) | 1.00 |
| P-DRUM-norm | 1.00 (0.00) | 0.82 (0.09) | 0.82 (0.25) | 0.99 (0.02) | 0.91 |
| P-DRUM-diff | 1.00 (0.00) | 0.97 (0.04) | 1.00 (0.00) | 1.00 (0.00) | 0.99 |
Appendix D When does P-DRUM outperform kNN and GMM?: an analysis based on principle component analysis (PCA)
In the in-domain setup, we observed that the P-DRUM-norm method performed the best across all methods; however, the baseline descriptor methods (kNN and GMM) also performed comparably on almost all datasets except HME21. Taking the dataset as an example (Figure 2), the descriptors in PC space for each atom in the train and test set were plotted, and each cluster in the plot represents the atoms of Ni or Al. The lower force error atoms in the figure represent atoms with lower force prediction error or low uncertainty. In the training set distribution, denser regions on the PCA plot correspond to lower force errors, i.e., smaller error magnitudes. This indicates that descriptor-based baseline methods such as GMM and kNN can achieve good performance without requiring prior knowledge of the force errors.
However, comparing to rMD17 elements and that contain at most 4 types of elements in each dataset, the HME21 dataset contains 37 different elements and a more diverse interaction between different elements. The landscape of prediction error is more complicated and difficult to learn without having the error information of the training set. Figure 3 shows the oxygen atoms, which is the most common atom in HME21 descriptors in PC space. We observed that in the train set, the circled area is the densest while having slightly higher force prediction error than the area on the right of the circle. kNN and GMM was unable to capture this and predicted the circled area as lowest uncertainty, while P-DRUM-norm method learned this from the forces prediction error during the training step.
Similar trends are observed in Figure 4, which only shows the calcium atoms in HME21. The top‑left region of the PCA plot is densely populated, yet exhibits relatively high force‑prediction errors. Both of our P-DRUM methods successfully capture this behavior, as the prediction error was explicitly incorporated into the training process. In contrast, the kNN and GMM approaches rely solely on descriptor information and therefore incorrectly assign low uncertainty to the same tail region in the top left. These results suggest that P-DRUM methods have strong potential for uncertainty prediction in more diverse and complex datasets, paving the way toward the development of universal potentials uncertainty estimation.
@@ Appendix D When does P-DRUM outperform kNN and GMM?: an analysis based on principle component analysis (PCA)
Justification: We propose P-DRUM and evaluate its performance to show the capability of our method.
@@
Justification: We share the training information and disclose training hyperparameters in Appendix B.
@@
Justification: We do not plan to release the code for the workshop submission version.
@@
Justification: We listed hyperparameters in Appendix B and describe training dataset in the main body. However, we also used dataset which is not open source but we believe we provided enough information for readers to understand the results.
@@
Justification: We described this information in Appendix B.
@@
Justification: We provided information of how dataset was generated in the experiment section.